OpenStudy (anonymous):

solve the following. (X-8) (X+7) = 5

OpenStudy (anonymous):

@UnkleRhaukus , help please?

OpenStudy (unklerhaukus):

expand the left hand side of the equation using foil \[(\overbrace{\overbrace{\color{orange}a +\underbrace{\color{cornflowerblue}b)(\color{seagreen}c}_{\text{Inside}}}^{\text{First}} +\color{brown}d}^{\text{Outside}}) =\overbrace{\color{orange}a\color{seagreen}c}^\text F +\overbrace{\color{orange}a\color{brown}d}^\text O +\underbrace{\color{cornflowerblue}b\color{seagreen}c}_\text I +\underbrace{\color{cornflowerblue}b\color{brown}d}_\text L\\ \qquad\quad \underbrace{\qquad\qquad}_{\text{Last}}\]

OpenStudy (anonymous):

How'd you do the colors?

OpenStudy (unklerhaukus):

majix

OpenStudy (anonymous):

(x^2+7x-8x-56)=5 (x^2-x-56)=5 I subtract 56 from the 5, right? Or vice versa?

OpenStudy (unklerhaukus):

x^2-x-56 = 5 is good work now take-away 5 from both sides

OpenStudy (anonymous):

x^2-x-61

OpenStudy (unklerhaukus):

yeah, x^2-x-61 = 0 now you could use completing the square method, or the quadratic formula

OpenStudy (unklerhaukus):

(the solutions very aren't nice, they have fractions and square-roots in them)

OpenStudy (anonymous):

Ooooooh nooooooo ! :/

hero (hero):

Just use the quadratic formula. It will save you from your fear.

OpenStudy (unklerhaukus):

A Quadratic Equation\[ax^2+bx+c=0\] The Quadratic Formula\[x_{1,2}=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] Factored for of the Quadratic Equation\[(x-x_1)(x-x_2)=0\] ___________ Start by identifying the coefficients a,b,c Compare \[x^2-x-61=0\] With \[ax^2+bx+c=0\] \[a=?\\ b=?\\ c=?\]

OpenStudy (unklerhaukus):

@cookiibabii93, can you find a,b,c ?

OpenStudy (anonymous):

a=x, b=x, c=-61

OpenStudy (unklerhaukus):

c = -61 is right but not a nor b

OpenStudy (anonymous):

a & b is 1

OpenStudy (unklerhaukus):

a =1 b = c= -61

OpenStudy (anonymous):

-1

OpenStudy (unklerhaukus):

thats better

OpenStudy (unklerhaukus):

plug these into the quadratic formula to get \[x_{1,2}=\frac{-(-1)\pm\sqrt{(-1)^2-4(1)(-61)}}{2(1)}\]

OpenStudy (anonymous):

\[\frac{ 1\pm \sqrt{1-244} }{ 2 }\]

OpenStudy (unklerhaukus):

careful the minus should be a plus

OpenStudy (anonymous):

Ooooh, there's where I messed up at, my answer was negative...I knew that was wrong

OpenStudy (anonymous):

\[\frac{ 1\pm \sqrt{245} }{ 2 }\]

OpenStudy (unklerhaukus):

very good, now you can simplify a little bit further what are factors of 245?

OpenStudy (anonymous):

\[7\sqrt{5}\]

OpenStudy (mrnood):

@unkleRhaukus Please tell how you did the colours and brackets and notes etc. in your earlier explanation...

OpenStudy (anonymous):

So you still didn't get the answer ???

OpenStudy (anonymous):

Nope...

OpenStudy (anonymous):

If multiplied, u gotta use FOIL (x-8) (x+7) = 5 x^2 + 7x -8x -56 = 5 x^2 - x -56 = 5 x^2 - x -61 = 0 now solve x using quadratic equation

OpenStudy (anonymous):

So u get 2 possible answers for x for this both are correct x = -7.33 x = 8.33

OpenStudy (anonymous):

awwww, thank yooooou !

OpenStudy (anonymous):

UnkleRhaukus clearly explained that .. lol

OpenStudy (anonymous):

And np if i understand something i'll help

OpenStudy (anonymous):

lol, yea, i just didnt feel like doing it lol